OpenStudy (anonymous):

compute the difference quotient of the given function Z(y)=3-8y-y^2

OpenStudy (anonymous):

same as your previous problem.... you'll need to compute: \(\large Z(y+h) \) for \(\large Z(y)=3-8y-y^2 \) then simplify the expression: \(\large \frac{Z(y+h)-Z(y)}{h} \)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would I plug this one in the same way as the other one?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

yeah the procedure is the same.... just a different function.... \(\large Z(y+h)=3-8(y+h)-(y+h)^2 \) simplify that...

OpenStudy (anonymous):

expand and simplify that expression then put that into the difference quotient.... sorry.... i gotta leave.... someone should be by to help u.... do you want me to tag someone i recommend?

OpenStudy (anonymous):

@Callisto , please help...

OpenStudy (anonymous):

ok thanks @dpalnc

OpenStudy (callisto):

Hello, what's going on here?! May I help?

OpenStudy (anonymous):

yesss... longer problems are harder for me to work out especially if I haven't yet experienced calculus yet...

OpenStudy (anonymous):

@callisto

OpenStudy (callisto):

Can you work out what Z(y+h) is?

OpenStudy (anonymous):

=3-8

OpenStudy (callisto):

Huh?! Z(y)=3-8y-y^2 Z(y+h) => replace y by y+h in Z(y)=3-8y-y^2.. So, what do you get?

OpenStudy (anonymous):

I don't know.... i"ll try to figure it out on my own thanks

OpenStudy (callisto):

Z(y+h) = 3-8(y+h) - (y+h)^2.. Since you have to replace y by y+h in Z(y) Then, you have to expand it. What do you get from the expansion?